A machine has a velocity ratio of 8 and efficiency of 85%. Calculate the mechanical advantage of the machine
1
Expert's answer
2020-05-08T16:06:41-0400

"efficiency=\dfrac{M.A.}{V.R.}"


"\dfrac{85}{100}=\dfrac{M.A.}{8}"


"M.A.=6.8"